4. When a stationary mass of gas was compressed without friction at constant

pressure its initial state of 0.4 m3 and 0.105 MPa was found to change to final state of 0.20 m3 and

0.105 MPa. There was a transfer of 42.5 kJ of heat from the gas during the process.

How much did the internal energy of the gas change ?


1
Expert's answer
2021-09-07T09:50:13-0400

"Q=\Delta U+A,"

"A=p\Delta V,"

"\Delta U=Q-p\Delta V,"

"\Delta U=42500-105000\cdot(0.4-0.2)=21.5~kJ."